# CLIP-Guided Backdoor Defense through Entropy-Based Poisoned Dataset Separation

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelandisbn: 978-1-4503-XXXX-X/18/06submissionid: #845ccs: Computing methodologies Machine learningccs: Security and privacy Software and application security
@@ 1. Introduction
We conducted extensive experiments to validate the effectiveness of CGD. With a runtime under 3 minutes, CGD can impressively reduce attack success rates (ASRs) to 0.2%, 0.6%, 1.0%, and 0.1% on average across 11 different attack categories for CIFAR-10, CIFAR-100, GTSRB, and Tiny-ImageNet, respectively, outperforming all existing defenses. Notably, we innovatively apply clean-data-based defenses (Liu et al., 2018; Wu and Wang, 2021) to poisoned data by filtering a clean subset with CGD. Results show that defenses using poisoned data can even surpass those using 5% clean data. Remarkably, even with a very weak CLIP (7.5% accuracy), our defense can still achieve a 0.7% ASR and only a 0.7% drop in clean accuracy. Additionally, we demonstrate that even when CLIP contains its own backdoor, CGD can still remove the backdoor from the victim model without transferring CLIP’s backdoor. Our contributions are threefold:
@@ 4.2.2. CLIP-Guided Neural Distillation
{tblr}
width = 1.0rowsep = 0.2pt, colsep = 0.1pt, colspec = Q[100]Q[100]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48]Q[48], cells = c, cell3-144 = bg = customred, cell3,4,8-116 = bg = customgreen, cell5-7,12-13,146 = bg = customred, cell3,5,7,8,12,138 = bg = customgreen, cell4,6,9-11,148 = bg = customred, cell3-5,7,8,1110 = bg = customgreen, cell6,9,10,12,13,1410 = bg = customred, cell3,4,8,1112 = bg = customgreen, cell5-7,9,10,12,13,1412 = bg = customred, cell3,6-1114 = bg = customgreen, cell4,5,12,13,1414 = bg = customred, cell3,4,6,7,8,1116 = bg = customgreen, cell5,9,10,12,13,1416 = bg = customred, cell3-1118 = bg = customgreen, cell12-1418 = bg = customred, cell3,4,6,7,10,1220 = bg = customgreen, cell5,8,9,11,13,1420 = bg = customred, cell3-1422 = bg = customgreen, cell11 = c=2, cell13 = c=2, cell15 = c=2, cell17 = c=2, cell19 = c=2, cell111 = c=2, cell113 = c=2, cell115 = c=2, cell117 = c=2, cell119 = c=2, cell121 = c=2, cell21 = c=2, cell31 = r=2, cell51 = r=4, cell91 = r=3, cell121 = r=2, cell141 = c=2, cell151 = c=4, cell161 = c=4, hline1,17 = -0.10em, hline2 = 3,5,7,9,11,13,15,17,19,21l, hline2 = 4,6,8,10,12,14,16,18,20,22r, hline3,5,9,12,14-15 = -0.05em, Defense $\rightarrow$ & No Defense ABL (Li et al., 2021b) DBD (Huang et al., 2021) ASD (Gao et al., 2023) D-BR (Chen et al., 2022) EP (Zheng et al., 2022) ReBack (Ma et al., 2024) PIPD (Chen et al., 2024) MSPC (Pal et al., 2024) CGD (ours)
@@ 5.1. Experimental Setups
{tblr}
width = 0.85rowsep = 0.2pt, colsep = 0.2pt, colspec = Q[202]Q[200]Q[200]Q[200]Q[200], cells = c, hline1,11 = -0.10em, hline2 = -0.05em, Method & (Prepare) Data Split Defense Total
@@ 5.2. Main Results
To evaluate the effectiveness of CGD, we report the CA and ASR results for eight defense methods against eleven attacks on CIFAR-10 in Table 1. CGD consistently achieves low ASRs ($\leq 1\%$) while maintaining high CAs, with minimal drops ($\leq 0.3\%$) across all defenses. Remarkably, CGD can identify and unlearn poisoned samples, improving CA over the ”No Defense” baseline by 3.4% for WaNet and 3.6% for IAB. In contrast, self-supervised defenses like DBD, ASD, and D-BR fail to resist clean-label attacks such as CTRL and SIG, while property-based defenses like ABL and EP are vulnerable to clean-image backdoors like GCB and FLIP. Additionally, CGD is highly efficient, requiring less than 3 minutes to split data and apply defenses after training the backdoored model (Table 2). While EP (Zheng et al., 2022) takes a similar amount of time, it cannot separate clean data. Thus, CGD stands out as the most efficient method for isolating clean data from poisoned samples while maintaining a high defense success rate.
@@
{tblr}
width = 0.85rowsep = 0.2pt, colsep = 0.2pt, colspec = Q[200]Q[87]Q[87]Q[87]Q[87]Q[87]Q[87]Q[87]Q[87], cells = c, cell4,6,7,10,11,12,13,149 = bg = customgreen, cell3,59 = bg = customred, cell10,12,13,145 = bg = customgreen, cell3-7,115 = bg = customred, cell6,10,133 = bg = customgreen, cell3,4,5,7,11,12,143 = bg = customred, cell3,5,6,7,11,127 = bg = customgreen, cell4,10,13,147 = bg = customred, cell12,4,6,8 = c=20.174, cell82,4,6,8 = c=20.174, hline1,15 = -0.10em, hline2 = 2,4,6,8l, hline2 = 3,5,7,9r, hline3,8,10 = -0.05em, hline9 = 2,4,6,8l, hline9 = 3,5,7,9r, Defense $\rightarrow$ & FP* (Liu et al., 2018) MCR* (Zhao et al., 2020) I-BAU* (Zeng et al., 2021) ANP* (Wu and Wang, 2021)
@@
Table 5.2.4 demonstrates the effectiveness of CGD across three additional datasets: CIFAR-100, GTSRB, and ImageNet. For each dataset, we evaluate all successful backdoor attacks (ASR $\geq$ 90%) and report the average results across attack categories. In all cases, CGD effectively eliminates backdoors from the models (with ASR reduced to $\leq$ 3%) while incurring only minimal clean accuracy (CA) reduction (average CA drop $\leq$ 0.3%). These results confirm the scalability and robustness of CGD across diverse datasets.
@@
{tblr}
width = 0.8rowsep = 0.2pt, colsep = 0.2pt, colspec = Q[250]Q[113]Q[113]Q[96]Q[119]Q[110]Q[110], cells = c, cell11 = r=2, cell12 = c=20.226, cell14 = c=20.215, cell16 = c=20.22, cell22 = c=20.226, cell24 = c=20.215, cell26 = c=20.22, cell91 = r=2, cell92 = c=20.226, cell94 = c=20.215, cell96 = c=20.22, cell102 = c=20.226, cell104 = c=20.215, cell106 = c=20.22, hline1,18 = -0.10em, hline3 = 2,4,6l, hline3 = 3,5,7r, hline11 = 2,4,6l, hline11 = 3,5,7r, hline4,9,12 = -0.05em, Dataset $\rightarrow$ & CIFAR-100  GTSRB  ImageNet
@@ 5.3.1. Loss Term Analysis
{tblr}
width = 1.0rowsep = 0.2pt, colsep = 0.2pt, colspec = Q[55]Q[55]Q[80]Q[60]Q[60]Q[65]Q[65]Q[74]Q[74]Q[71]Q[71]Q[60]Q[60], cells = c, cell11 = c=3, cell14 = c=2, cell16 = c=2, cell18 = c=2, cell110 = c=2, cell112 = c=2, vline4,12 = 1-90.05em, hline1,10 = -0.1em, hline3,9 = -0.05em, cell3-45,7,9,11,13 = bg=customred, cell55 = bg=customgreen, cell57,9,11,13 = bg=customred, cell65,7,9,11,13 = bg=customred, cell75,9,11 = bg=customgreen, cell77,13 = bg=customred, cell85,9,11,13 = bg=customgreen, cell87 = bg=customred, cell95,7,9,11,13 = bg=customgreen, Case Classic Dynamic Clean-label Clean-lmage Average
@@
{tblr}
width = 1.0rowsep = 0.2pt, colsep = 0.2pt, colspec = Q[260]Q[138]Q[138]Q[160]Q[160], cells = c, cell12 = c=2, cell14 = c=2, hline2 = 2,4l, hline2 = 3,5r, hline1,8 = -0.1em, hline3,7 = -0.05em, Model→ SigLIP ImageBind
@@ 5.3.3. Robustness to Different Language-Image Pretraining
Although we mainly use CLIP (Radford et al., 2021) to illustrate CGD’s effectiveness, CGD can use other language-image pretraining models to replace CLIP. As shown in Table 6, we consider SigLIP (Zhai et al., 2023) and ImageBind (Girdhar et al., 2023) as two alternative models. The result suggests that both achieve low ASR ($\leq 5\%$) while maintaining good CA ($\geq 90\%$).
Figure 6. CGD’s defense performance on low-accuracy CLIP models. CGD successfully defends against all attacks (ASR $\leq 5\%$) even when CLIP’s zero-shot accuracy is as low as 10%.
@@ 5.4.1. Weaker CLIP
To further validate this assumption in real and diverse vision datasets, we carry out very extensive experiments on 17 different datasets, including large-scale datasets like SUN397 and full ImageNet. On these datasets, we tested three representative attacks (Blend, BPP, CTRL) but excluded clean-image backdoors due to either low ASR on large datasets (FLIP) or high computational cost (GCB). As shown in Table 7, CGD successfully defends all attacks except on SVHN for BPP. This failure is because CLIP’s zero-shot accuracy on SVHN is only 13.4%, making it thoroughly ineffective for a 10-class prediction dataset. Nonetheless, CGD still defends other SVHN attacks (e.g., Blend, CTRL) via the suspicious model’s entropy. Regardless of whether the dataset is large-scale or has low CLIP accuracy (e.g., GTSRB, DTD, and MNIST), CGD remains effective, showing its strong generalizability.
@@
{tblr}
width = 1.0rowsep = 0.2pt, colsep = 0.8pt, colspec = Q[221]Q[92]Q[83]Q[83]Q[83]Q[83]Q[83]Q[83]Q[83]Q[83], cells = c, cell11 = r=2, cell12 = r=3, cell13 = c=2,r=2, cell15 = c=6, cell25 = c=2, cell27 = c=2, cell29 = c=2, cell43 = bg=customblue17, cell44 = bg=customblue0, cell53 = bg=customblue16, cell54 = bg=customblue17, cell63 = bg=customblue13, cell64 = bg=customblue0, cell73 = bg=customblue11, cell74 = bg=customblue7, cell83 = bg=customblue11, cell84 = bg=customblue5, cell93 = bg=customblue10, cell94 = bg=customblue0, cell103 = bg=customblue8, cell104 = bg=customblue9, cell113 = bg=customblue8, cell114 = bg=customblue9, cell123 = bg=customblue7, cell124 = bg=customblue7, cell133 = bg=customblue7, cell134 = bg=customblue5, cell143 = bg=customblue7, cell144 = bg=customblue5, cell153 = bg=customblue6, cell154 = bg=customblue2, cell163 = bg=customblue3, cell164 = bg=customblue1, cell173 = bg=customblue2, cell174 = bg=customblue1, cell183 = bg=customblue2, cell184 = bg=customblue5, cell193 = bg=customblue2, cell194 = bg=customblue1, cell203 = bg=customblue0, cell204 = bg=customblue0, cell4-206 = bg = customgreen, cell5-208 = bg = customgreen, cell48 = bg = customred, cell4-2010 = bg = customgreen, hline3 = 3,5,7,9,11l, hline3 = 4,6,8,10,12r, hline2 = 5-11, hline1,21 = -0.1em, hline4 = -0.05em, Task→ Data
Size
Classify
Accuracy
CGD (Ours)
@@ 5.5. Resistance to Potential Adaptive Attacks
Table 9 demonstrates that our CGD defense effectively mitigates both adaptive attacks. Against Adaptive-SIG, CGD reduces ASRs to $\leq 2.4\%$. As shown in Fig. 7, while some triggered samples evade detection during the splitting stage, they retain minimal trigger information, with strongly triggered samples all identified and unlearned. For the FMB attack, ASR is also reduced to $\leq 1.2\%$ across both one-to-one and all-to-one settings. This occurs because, although the CLIP model may confuse mixed-feature samples, it only misclassifies them within the two mixed classes, not the target (third) class. Consequently, the poisoned target label also yields a high cross-entropy score, making the attack ineffective against our defense.
@@
{tblr}
width = 0.8rowsep = 0.8pt, colsep = 0.8pt, colspec = Q[250]Q[250]Q[110]Q[110]Q[110]Q[110], cells = c, cell11 = c=2, cell13 = c=2, cell15 = c=2, cell21 = c=2, cell31 = r=3, cell61 = r=2, hline1,8 = -0.1em, hline2 = 3,5l, hline2 = 4,6r, hline3,6 = -0.05em, Defense → No Defense CGD (Ours)
@@ A.1. Detection of Label-Poisoned Samples with CLIP
Consider a dataset $D=\{(x_{i},y_{i})\}_{i=1}^{N}$, where a fraction $pr$ ($10<pr<$ 1) of samples are poisoned with incorrect labels. The clean subset is $C$ ($|C|=(1-pr)N$), and the poisoned subset is $P$ ($|P|=prN$). For each sample, CLIP computes a cross-entropy score $\mathcal{S}_{i}^{\text{CLIP}}=-\log p(y_{i}\mid x_{i})$, where $p(y_{i}\mid x_{i})$ is the predicted probability of the assigned label $y_{i}$. We rank samples by $\mathcal{S}_{i}^{\text{CLIP}}$ in descending order and flag the top $k=prN$ as potentially poisoned.
@@ D.1. Effectiveness Across Poison Rates
{tblr}
width = rowsep = 0.8pt, colsep = 0.8pt, colspec = Q[200]Q[83]Q[102]Q[88]Q[85]Q[92]Q[92]Q[88]Q[87], cells = c, cell12 = c=40.358, cell16 = c=40.359, cell22 = c=20.185, cell24 = c=20.175, cell26 = c=20.184, cell28 = c=20.175, cell21 = r=2, hline1,16 = -0.10em, hline2,4,15 = -0.05em, hline3 = 2-90.03em, Poison Rate 1% 10%
@@
{tblr}
width = rowsep = 0.8pt, colsep = 0.8pt, colspec = Q[200]Q[83]Q[102]Q[88]Q[85]Q[92]Q[92]Q[88]Q[87], cells = c, cell12 = c=40.358, cell16 = c=40.359, cell22 = c=20.185, cell24 = c=20.175, cell26 = c=20.184, cell28 = c=20.175, cell21 = r=2, hline1,16 = -0.08em, hline2,4,15 = -0.05em, hline3 = 2-90.03em, Architecture VGG16 MobileNet V2
@@ Appendix E More Results in Ablation Study
{tblr}
width = 1.0rowsep = 0.5pt, colsep = 0.5pt, colspec = Q[150]Q[60]Q[60]Q[60]Q[60]Q[80]Q[80]Q[80]Q[80]Q[60]Q[60], cells = c, cell12 = c=2, cell14 = c=2, cell16 = c=2, cell18 = c=2, cell110 = c=2, cell3-53 = bg=customgreen, cell3,55 = bg=customgreen, cell45 = bg=customred, cell4,57 = bg=customgreen, cell37 = bg=customred, cell3,59 = bg=customgreen, cell49 = bg=customred, cell3,511 = bg=customgreen, cell411 = bg=customred, vline2,10 = 1-50.05em, hline1,6 = -0.1em, hline3 = -0.05em, Category → Classic Dynamic Clean-label Clean-image Average
@@ Appendix F Use CLIP Variants in Domain-Specific Tasks
In other domain-specific environments like medical imaging, CLIP models such as PubMed (Eslami et al., 2023) (62.7% accuracy on PCAM) achieved an ASR of 6.1%, as shown in Table 13. These results demonstrate that with advanced domain-specific CLIP variants (e.g. PubMedCLIP (Eslami et al., 2023) in Medicine, BioCLIP (Stevens et al., 2024) in Biology, GeomCLIP (Xiao et al., 2024) in Chemistry), defenders can select high-performing CLIP variants on the poisoned dataset to ensure CGD’s effectiveness. CLIP’s public availability and ongoing advancements in domain-specific models further reduce feasibility concerns.
@@
{tblr}
width = 0.8rowsep = 0.2pt, colsep = 0.8pt, colspec = Q[238]Q[254]Q[192]Q[113]Q[113], cells = c, cell12 = r=2, cell13 = r=2, cell14 = c=2, cell31 = r=3, cell61 = r=3, hline1,9 = -0.1em, hline3,6 = -0.05em, Task→ CLIP Weight CLIP Acc. BPP Attack
@@
{tblr}
width = 1.0rowsep = 0.2pt, colsep = 0.1pt, colspec = Q[94]Q[71]Q[46]Q[46]Q[46]Q[46]Q[46]Q[46]Q[46]Q[54]Q[46]Q[46]Q[46]Q[46]Q[48]Q[48]Q[46]Q[46], cells = c, cell11 = c=2, cell13 = c=8, cell111 = c=8, cell21 = c=2, cell23 = c=2, cell25 = c=2, cell27 = c=2, cell29 = c=2, cell211 = c=2, cell213 = c=2, cell215 = c=2, cell217 = c=2, cell31 = c=2, cell41 = r=2, cell61 = r=4, cell101 = r=3, cell131 = r=2, cell151 = c=2, cell4,6,9,134 = bg = customgreen, cell5,7,8,10,11,12,14,154 = bg = customred, cell4,6-10,13-156 = bg = customgreen, cell5,11,126 = bg = customred, cell4,6,9,11-158 = bg = customgreen, cell5,7,8,108 = bg = customred, cell4,6,7,1310 = bg = customgreen, cell5,8-12,14,1510 = bg = customred, cell4,6,8,9,12-1412 = bg = customgreen, cell5,7,10,11,1512 = bg = customred, cell4,6-1514 = bg = customgreen, cell514 = bg = customred, cell4,6-9,11-13,1516 = bg = customgreen, cell5,10,1416 = bg = customred, cell4-13,1518 = bg = customgreen, cell1418 = bg = customred, vline3,11 = -0.05em, hline1,16 = -0.1em, hline4,6,10,13,15 = -0.05em, hline3 = 3,5,7,9,11,13,15,17,19l, hline3 = 4,6,8,10,12,14,16,18,20r, Requirement → 5% Clean Training Data 100% Poisoned Training Data
@@
{tblr}
width = 1.0rowsep = 0.2pt, colsep = 0.1pt, colspec = Q[94]Q[71]Q[46]Q[46]Q[46]Q[46]Q[46]Q[46]Q[46]Q[54]Q[46]Q[46]Q[46]Q[46]Q[48]Q[48]Q[46]Q[46], cells = c, cell11 = c=2, cell13 = c=8, cell111 = c=8, cell21 = c=2, cell23 = c=2, cell25 = c=2, cell27 = c=2, cell29 = c=2, cell211 = c=2, cell213 = c=2, cell215 = c=2, cell217 = c=2, cell31 = c=2, cell41 = r=2, cell61 = r=4, cell101 = r=3, cell131 = r=2, cell151 = c=2, cell4,6,9,134 = bg = customgreen, cell5,7,8,10,11,12,14,154 = bg = customred, cell4,6-10,13-156 = bg = customgreen, cell5,11,126 = bg = customred, cell4,6,9,11-158 = bg = customgreen, cell5,7,8,108 = bg = customred, cell4,6,7,1310 = bg = customgreen, cell5,8-12,14,1510 = bg = customred, cell4,6,8,9,1312 = bg = customgreen, cell5,7,10-12,14,1512 = bg = customred, cell4,6-11,13-1514 = bg = customgreen, cell5,1214 = bg = customred, cell4,7-13,1516 = bg = customgreen, cell5,6,1416 = bg = customred, cell6-818 = bg = customgreen, cell4,5,9-1518 = bg = customred, vline3,11 = -0.05em, hline1,16 = -0.1em, hline4,6,10,13,15 = -0.05em, hline3 = 3,5,7,9,11,13,15,17,19l, hline3 = 4,6,8,10,12,14,16,18,20r, Requirement → 5% Clean Training Data 10% Poisoned Training Data

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3754860isbn: 979-8-4007-2035-2/2025/10submissionid: #845ccs: Computing methodologies Machine learningccs: Security and privacy Software and application security
@@ 1. Introduction
We conducted extensive experiments to validate the effectiveness of CGD. With a runtime under 3 minutes, CGD can impressively reduce attack success rates (ASRs) to 0.2%, 0.6%, 1.0%, and 0.1% on average across 11 different attack categories for CIFAR-10, CIFAR-100, GTSRB, and Tiny-ImageNet, respectively, outperforming all existing defenses. Notably, we innovatively apply clean-data-based defenses (Liu et al., 2018; Wu and Wang, 2021) to poisoned data by filtering a clean subset with CGD. Results show that defenses using poisoned data can even surpass those using 5% clean data. Remarkably, even with a very weak CLIP (7.5% accuracy), our defense can still achieve a 0.7% ASR and only a 0.7% drop in clean accuracy. Additionally, we demonstrate that even when CLIP contains its own backdoor, CGD can still remove the backdoor from the victim model without transferring CLIP’s backdoor.
Our contributions are threefold:
@@ 4.2.2. CLIP-Guided Neural Distillation
Defense $\rightarrow$ No Defense ABL (Li et al., 2021b) DBD (Huang et al., 2021) ASD (Gao et al., 2023) D-BR (Chen et al., 2022) EP (Zheng et al., 2022) ReBack (Ma et al., 2024) PIPD (Chen et al., 2024) MSPC (Pal et al., 2024) CGD (ours)
@@ 5.1. Experimental Setups
Method (Prepare) Data Split Defense Total
@@ 5.2. Main Results
[…]
@@
Defense $\rightarrow$ FP* (Liu et al., 2018) MCR* (Zhao et al., 2020) I-BAU* (Zeng et al., 2021) ANP* (Wu and Wang, 2021)
@@
Table 4 demonstrates the effectiveness of CGD across three additional datasets: CIFAR-100, GTSRB, and ImageNet. For each dataset, we evaluate all successful backdoor attacks (ASR $\geq$ 90%) and report the average results across attack categories. In all cases, CGD effectively eliminates backdoors from the models (with ASR reduced to $\leq$ 3%) while incurring only minimal clean accuracy (CA) reduction (average CA drop $\leq$ 0.3%). These results confirm the scalability and robustness of CGD across diverse datasets.
@@
Dataset $\rightarrow$ CIFAR-100 GTSRB ImageNet
@@ 5.3.1. Loss Term Analysis
Case Classic Dynamic Clean-label Clean-lmage Average
@@
Model→ SigLIP ImageBind
@@ 5.3.3. Robustness to Different Language-Image Pretraining
[…]
@@ 5.4.1. Weaker CLIP
To further validate this assumption in real and diverse vision datasets, we carry out very extensive experiments on 17 different datasets, including large-scale datasets like SUN397 and full ImageNet. On these datasets, we tested three representative attacks (Blend, BPP, CTRL) but excluded clean-image backdoors due to either low ASR on large datasets (FLIP) or high computational cost (GCB). As shown in Table 5.4.1, CGD successfully defends all attacks except on SVHN for BPP. This failure is because CLIP’s zero-shot accuracy on SVHN is only 13.4%, making it thoroughly ineffective for a 10-class prediction dataset. Nonetheless, CGD still defends other SVHN attacks (e.g., Blend, CTRL) via the suspicious model’s entropy. Regardless of whether the dataset is large-scale or has low CLIP accuracy (e.g., GTSRB, DTD, and MNIST), CGD remains effective, showing its strong generalizability.
@@
Task→ Data
Size Classify
Accuracy CGD (Ours)
@@ 5.5. Resistance to Potential Adaptive Attacks
[…]
@@
Defense → No Defense CGD (Ours)
@@ A.1. Detection of Label-Poisoned Samples with CLIP
[…]
@@ D.1. Effectiveness Across Poison Rates
Poison Rate 1% 10%
@@
Architecture VGG16 MobileNet V2
@@ Appendix E More Results in Ablation Study
Category → Classic Dynamic Clean-label Clean-image Average
@@ Appendix F Use CLIP Variants in Domain-Specific Tasks
In other domain-specific environments like medical imaging, CLIP models such as PubMed (Eslami et al., 2023) (62.7% accuracy on PCAM) achieved an ASR of 6.1%, as shown in Table F. These results demonstrate that with advanced domain-specific CLIP variants (e.g. PubMedCLIP (Eslami et al., 2023) in Medicine, BioCLIP (Stevens et al., 2024) in Biology, GeomCLIP (Xiao et al., 2024) in Chemistry), defenders can select high-performing CLIP variants on the poisoned dataset to ensure CGD’s effectiveness. CLIP’s public availability and ongoing advancements in domain-specific models further reduce feasibility concerns.
@@
Task→ CLIP Weight CLIP Acc. BPP Attack
@@
Requirement → 5% Clean Training Data 100% Poisoned Training Data
@@
Requirement → 5% Clean Training Data 10% Poisoned Training Data